OpenStudy (anonymous):

5. What is a cubic polynomial function in standard form with zeros 1, –2, and 2?

OpenStudy (amistre64):

subtract x from the given zeros and multiply them all together

OpenStudy (amistre64):

1, –2, 2 1-x, –2-x, 2-x f(x) = (1-x) (–2-x) (2-x)

OpenStudy (anonymous):

Choices given. f(x)= x3 + x2 – 3x + 4 f(x)= x3 + x2 – 4x – 2 f(x)= x3 + x2 + 4x + 4 f(x)= x3 – x2 – 4x + 4

OpenStudy (amistre64):

then you would need to pass a negative thru my results to get one of those

OpenStudy (sirm3d):

subtract each root from its own x, then multiply the binomials formed.